The function: f(x) = |sin(x) - 1| / (sin(x) - 1)
can be redefined to make it continuous at x= pi/2

Answer by ikleyn(52781)

.


        As I read your post, it is unclear to me, is it a statement or a question.

        But, taking into account, that half of the posts at this forum are mathematically
        and grammatically incorrect and require to be fixed,
        I will assume that it is a question of the type "True or False".


The given function is identically equal to -1 everywhere in vicinity of x = pi/2
and is not defined at x = pi/2.

Therefore, if to define it as -1 at x= pi/2, you will get a continuous function at x= pi/2.